a property owner wants to fence a garden plot adjacent to a road. the fencing next to the road must be sturdier and costs $5 per foot and the other fencing just $3 per foot. the garden has to have an area of 1200 square feet. find the garden dimensions that he can fence-in he spends $600
1
Expert's answer
2012-11-12T11:42:12-0500
Let x and y are the dimensions 0f the garden and xis the side next to the road. The area of the garden is x*y=1200. He spends:
5x+3x+2*3y=8x+6y=600.
Then we have the system of the equations:
x*y=1200
8x+6y=600
y=(600-8x)/6
600x-8x^2=7200
x^2-75x+900=0
x1=15, y1=80
x2=60, y2=20
Dimensions of the garden are (15,80) foot or (60,20) feet.